OpenStudy (anonymous):

\[log(4x+5) - log(3x+2)\]

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

\[log(a) - log(b) = log\left(\frac{a}{b}\right)\]

OpenStudy (anonymous):

log[(4x+5)/(3x+2)]

OpenStudy (anonymous):

\[\log { \frac{4x + 5 }{3x + 2 }} \]

OpenStudy (anonymous):

can we find the value of x by the given equations ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hmm is it equal to zero

OpenStudy (anonymous):

k thanks very much Ishaan bhaiya : Sajan bhaiya and Dalvoron

OpenStudy (anonymous):

If there is no equals sign, we cannot find the value.

OpenStudy (anonymous):

ya if it has an equal sign

OpenStudy (anonymous):

\[log(4x+5) - log(3x+2)=0\] ?

OpenStudy (anonymous):

k \[log(4x+5) = log(3x+2)\]

OpenStudy (anonymous):

\[log(4x+5)=log(3x+2)\]\[\rightarrow 4x+5=3x+2\]

OpenStudy (anonymous):

can we solve it now ? how if yes and no ?

OpenStudy (anonymous):

\[\log { \frac{4x + 5 }{3x + 2 }}= 0 \] \[ \log { \frac{4x + 5 }{3x + 2 }}= \log 1 \] \[ 4x + 5 = 3x + 2\] \[x = -3\]

OpenStudy (anonymous):

dalvoron u cut out log ? can we cut log

OpenStudy (anonymous):

Yes, if \(log(a)=log(b)\), then a = b.

OpenStudy (anonymous):

thanks ishaan bhaiya and dalvoron again please comment out there in my next question ( comment only OK )